Reported magazine newsstand sales were used to predict audited sales for 10 magazines. For that data, SSR = 132,184.22 and SSWill rate, thank you in advance.
Answer #1

Solution:

n = 10 , SSR = 132,184.22 , SST = 144,410.99 , SSE = SST - SSR = 144,410.99-132,184.22 = 12226.77

r^2 = SSR/SST = 132,184.22/144,410.99 = 0.9153

a.

r^2 = 0.9153

It means that 91.53 % of the variation in audited newsstand sales is be explained by the variation in reported newsstand sales.

b.

Sxy = √SSE/n-2 = √12226.77/8 = 39.0941

Sxy = 39.0941

c.

  • A. It is very useful for predicting audited sales because the coefficient of determination is close to 1 and the standard error of estimate is fairly small.